One adult and 10 children are in an elevator. If the adults weight is 4 times the average weight of the children, the the adults weight is what fraction of the total weight of the 11 people in the elevator?

Answer by stanbon(75887)

One adult and 10 children are in an elevator. If the adults weight is 4 times the average weight of the children, the adult's weight is what fraction of the total weight of the 11 people in the elevator?
----------
Let the average wt. of each child be "c".
Wt of the one adult is 4c
------
Total wt in the elevator is 10*c+4c = 14c
-----
Ans:: The adult's wt is 4c/14c = 2/7
-------------
Cheers,
Stan H.
------------